What decimal number does the bit pattern 10101100 represent if it is a:

  • [1 pts] unsigned integer?
  • [1 pts] sign-magnitude integer?
  • [1 pts] two's complement integer?

What decimal number does the bit pattern 01010011 represent if it is a:

  • [1 pts] unsigned integer?
  • [1 pts] sign-magnitude integer?
  • [1 pts] two's complement integer?
Answer #1

1. Given bit pattern : 10101100

a) if it is unsigned integer : 1 * 27 + 1 * 25 + 1 * 23 + 1 * 22

= 128 + 32 + 8 + 4 = 172

b) If it is a sign-magnitude integer : 10101100 -- MSB is 1. So it is a negative integer

0101100 = 1 * 25 + 1 * 23 + 1 * 22 = 32 + 8 + 4 = 46

= -46

c) If it is two's complement integer :  10101100 -- MSB is 1. So it is a negative number.

-- first invert the bits ====> 01010011

-- Now add 1 to it. =====> 01010011

+ 1

--------------------

01010100

magnitude of 01010100 = 1 * 26+ 1 * 24  + 1 * 22 = 64 + 16 + 4 = 84

So = -84.

2. Given bit pattern 01010011

a) if it is unsigned integer : 1 * 26+ 1 * 24  + 1 * 21 + 1 * 20 = 64 + 16 + 2 + 1 = 83

b) If it is a sign-magnitude integer : 01010011 - MSB is 0. So it is a positive number.

So the integer = 83

c) If it is two's complement integer :  01010011 -- MSB is 0. So it is a Positive number.

So the integer = 83, because for positive 2's complement we dont need to do anything.
Answer #2

For the bit pattern 10101100:

  1. Unsigned integer:

    • Think of it as a plain binary number.

    • Calculation: 1×2⁷ + 0×2⁶ + 1×2⁵ + 0×2⁴ + 1×2³ + 1×2² + 0×2¹ + 0×2⁰ = 128 + 0 + 32 + 0 + 8 + 4 + 0 + 0 = 172.


  2. Sign-magnitude integer:

    • The first bit is the sign (1 = negative), the rest is the magnitude.

    • Magnitude: 0101100 = 0×2⁶ + 1×2⁵ + 0×2⁴ + 1×2³ + 1×2² + 0×2¹ + 0×2⁰ = 0 + 32 + 0 + 8 + 4 + 0 + 0 = 44.

    • Final value: -44.


  3. Two's complement integer:

    • The first bit is negative, so invert the bits and add 1 to find the positive equivalent.

    • Original: 10101100

    • Invert: 01010011

    • Add 1: 01010100 = 84

    • Final value: -84.


For the bit pattern 01010011:


  1. Unsigned integer:

    • Plain binary number.

    • Calculation: 0×2⁷ + 1×2⁶ + 0×2⁵ + 1×2⁴ + 0×2³ + 0×2² + 1×2¹ + 1×2⁰ = 0 + 64 + 0 + 16 + 0 + 0 + 2 + 1 = 83.


  2. Sign-magnitude integer:

    • First bit is 0 = positive, the rest is the magnitude.

    • Magnitude is the same as above: 83.

    • Final value: +83.


  3. Two's complement integer:

    • First bit is 0, so it's the same as the unsigned value.

    • Final value: +83.


Summary :-

  • 10101100: Unsigned = 172, Sign-magnitude = -44, Two's complement = -84.

  • 01010011: Unsigned = 83, Sign-magnitude = +83, Two's complement = +83.


answered by: Harshwardhan kunal